If a spinning object with a radius of 2 m, a mass of 7 kg, and an angular momentum of 21kg*m^2/s, breaks into two equal pieces, what will be the final velocity of each piece? a. 4.5 m/s b. 3 m/s c. 1.5 m/s d. 0 m/s
Answer #1

Initial angular momentum = 21 kg*m2/s
Now we know that the spinning object is break into two equal pieces therefore
Final angular momentum of the system = 2(mVr)
where m is mass of new piece = 7/2 = 3.5 kg
V is speed of each piece
r is radius = 2 m
Now applying conservation of angular momentum
21 = 2(3.5*V*2)
V = 1.5 m/s
Hence the correct option will be C.